OpenStudy (anonymous):

how to simplify sqrt(x^2-16)

OpenStudy (anonymous):

didn't even think of that haha, thank you

OpenStudy (anonymous):

square the square root

OpenStudy (anonymous):

so i have sqrt((x+4)(x-4)) what can i do next?

OpenStudy (anonymous):

you cannot simplify it

OpenStudy (anonymous):

i thought you can square the square root

OpenStudy (anonymous):

you can rewrite the inside in factored form, but that is in no way simplified and no, you cannot square it that is like saying i can simplify \(2\) by writing \(4\)

OpenStudy (anonymous):

is this the actual problem, or is it part of another problem?

OpenStudy (anonymous):

this is part of a larger problem.. here is the whole problem: f(x)=sqrt(x^2-16) g(x)=x^2/(x^2+1) find (f+g)(x)

OpenStudy (anonymous):

then just stick a plus sign between them

OpenStudy (anonymous):

alright thanks, its online hw so i'll check if that's as far as it'll go